# Evaluating the Limits of QAOA Parameter Transfer at High-Rounds on Sparse Ising Models With Geometrically Local Cubic Terms

### v1

@@ III.4 Validating Large Scale QAOA Parameter Transfer with PEPS Simulations
Figure 10 plots the Hamiltonian expectation value as a function of $p$, for the complete set of $12$ trained sets of QAOA angles, using PEPS with a bond dimension of $D=64$, for the $127$, $133$, and $156$ qubit Ising model instances. The goal here is to validate whether the parameter transfer continues to work at larger problem sizes. Indeed, we see that for some of the QAOA angles, parameter transfer works remarkably well – allowing convergence to nearly the true ground-state. However, similarly to the smaller scale statevector simulations of Figures 4, 5 and 6, we see that for some sets of parameters, there exist non-monotonic fluctuations, dips, and in one case there is a single set of QAOA angles which results in a significantly worse performance (note that this angle is shown explicitly as the highly fluctuating schedule shown in the 5th row from the top of Figure 1). This result shows a single case where non-smooth QAOA angles can perform well on the small test case problem instance, but then fail to perform well when those parameter are transferred up to larger problem sizes – we hypothesize that this is likely rare artifact associated with having converged to close the ground-state energy, although additional investigation is required on this point.
In summary, Figure 10 shows that QAOA parameter transfer can work well across instances which have a parameter-transfer Ising model size difference (between the trained instance and the tested instance) of up to $140$ discrete variables, and in this case allows us to get near to the true ground-state, assuming a noise-less quantum computation (approximated by PEPS with a bond dimension of $D=64$).
Figure 11: Validating the QAOA angle quality, trained on the set of $12$ distinct $27$-qubit Ising model instances, applied to the $156$-qubit (left), $133$-qubit (middle), and $127$-qubit (right) instances, using $D=64$ PEPS simulations with belief propagation for approximate tensor network contraction. Ground state energy not marked because these expectation values $\leq 7$ do not reach near the ground state, as opposed to Figure 10 where the mean expectation value does approach the true ground state energy.
@@
Figure 13: Ground-state energy QAOA sampling rate (probability of sampling the ground state) from $D=64$ and $D=32$ PEPS simulations, across the ensemble of $16$-qubit trained angles, on the $156$-qubit Ising model (left), $133$-qubit Ising model (middle) and the $127$-qubit Ising model (right). The ground-state sampling proportion is estimated using $1000$ samples, per parameter, (10,00010{,}000 in the case of the $133$-qubit case, to reduce the effect of shot noise due to the low ground-state energy sampling rate) and each sub-plot y-axis is different for each Ising model. Interestingly, the $133$-qubit instance ground-state is rarely sampled, in comparison to the other two models.
Figure 11 shows similar trends of monotonic expectation value improvement as a function of $p$ from the $27$-qubit instance trained angles, as we would want from good QAOA evolution – but with one exception of the entirely positive coefficient Ising model having a non-monotonic dip in performance. Comparing Figure 10 and Figure 11, there is little difference between the angles that were trained either the slightly smaller ($16$ qubit) or larger ($27$ qubit) instances.
@@
Interestingly, one of the primary conclusions from all of these QAOA parameter transfer calculations is that overall, coefficient choice does not significantly impact the transfer of parameters (the only clear exception to this is that parameters from the positive coefficient case do not monotonically perform well on the other instances). This suggests that primarily, the Hamiltonian connectivity structure impacts the transfer of parameters much more significantly than the cost function coefficient distribution.
@@
Figure 12 plots the relative expectation value differences from PEPS simulations at different bond dimensions run on the largest Ising model we consider in this study ($156$ spins). This shows that the tensor network simulations clearly converge at small $p$ (typically $p\leq 7$) and typically fail to converge as strongly at larger $p$, with slower convergence occurring in the intermediate values of $p$ typically $p=10$ to =20=20. Moreover, the trend of PEPS convergence is quite distinct for each set of fixed QAOA angles. This empirical convergence evidence reflects what we would expect from highly optimized QAOA evolution [96], similarly to quantum annealing, where in the middle of the evolution the system has high entanglement entropy, thus making it harder for the PEPS simulation to converge.
@@
Figure 18 in Appendix D shows the convergence properties of the remaining set of trained QAOA angles, run on the single $156$-qubit instance. Interestingly, this lack of full convergence at intermediate p values (typically this is approximately between $p=10$ and $p=30$) suggests that those specific QAOA angles result in circuits which are harder to approximate classically using tensor network methods. This convergence trend manifests across most of the angles, with the one exception of the all negative-coefficient angle set shown in Figure 18. However, the larger $p$ simulations, which use slightly different QAOA angles, are then closer to convergence. This suggests that there do exist particular QAOA parameter combinations for these circuits which are hard to classically simulate, however, unexpectedly the optimized angles, which typically approach the true ground-state energy, are easier to approximate classically. This could be related to the properties of the ground-states themselves which are approached in the high-$p$ regime, which can be expressed as a linear combination of many product states (i.e., due to high degeneracy of the Ising models).
@@
Convergence of the PEPS simulations would resemble a clear progression downward on the y-axis of logarithm expectation value difference in Figure 12. Eventually at very large bond dimension, convergence results in the difference between bond dimensions (which are increasing as powers of 2) which have zero difference between each other. This effect is really only seen very strongly at low-p depth in Figure 12, and in some cases we see a strong lack of convergence at sufficiently high p, indicating those expectation values have higher uncertainty, even at $D=128$.
Lastly, a natural question in regards to quantum computation is the effect of sampling – for example Figure 10 showed only the mean expectation value from the PEPS simulation, namely because all quantum computation requires measuring the state of the qubits at the end of the computation.
@@
Figure 13 shows sampling from the PEPS simulation at $D=64$ using $1000$ samples (for each QAOA angle, at each $p$), where we quantify the sampled ground-state energy proportion. Interestingly, this shows that some of the parameter transfer angles work very well for ground-state sampling, typically well exceeding a proportion of $0.1$, while others do not perform well. This suggests in general for parameter transfer, using an ensemble of distinct QAOA angles is a good technique. However, the $133$-qubit instance ground-state energy is very rarely sampled. The gap between the best-performing angle and the ground-state energy of the $133$-qubit in Figure 10 is larger than the gap for the other two instances, and therefore the low ground-state sampling rate in Figure 13 is consistent with the noiseless expectation values. This suggests that the low ground state (GS) sampling rate is an instance-dependent result, and the performance of these fixed angles likely has high variability across different random-coefficient problem instances. Figure 13 reports only the QAOA angles originating from the training on $16$-qubit instances since the $27$-qubit instances were only trained up to $p=7$, and in that QAOA depth regime we verified with the PEPS simulations that GS energy sampling rate (using $1000$ samples per $p$) is zero.
@@ III.5 Validating QAOA PEPS Simulations with MPS Simulations
The MPS simulations are performed with $\chi=2^{k}$, for $k=7,\ldots,11$ and shown in Figure 14 for the $156$ qubit QAOA circuits. Panels (a) and (b) display the error in the energy $\Delta E$ as a function of $p$ for several values of MPS bond dimension $\chi$. Here, the error $\Delta E$ quantifies the difference between our most accurate PEPS energy approximation and the MPS approximation computed with a given $\chi$. It is defined as:
@@ IV Discussion and Conclusion
However, parameter transfer does have limitations. Concretely, the following limitations exist for QAOA parameter transfer, primarily for the task of training on smaller problem instances and applying those parameters to larger problems. The most direct limitation is that once the QAOA training saturates the maximum approximation ratio of $1$, meaning no further improvement is possible, then the training can not supply good angles past that saturation point (for larger $p$) to use on any larger problem instances. This saturation is a problem specifically when we generally expect higher QAOA depth simulations to be required, to reach a certain threshold of accuracy (such as a ground-state sampling rate), as problem size increases. This limit necessitates the development of other techniques to be able to efficiently compute good QAOA angles, such as extrapolating to higher QAOA depths based on the already learned angles.
@@ Appendix A Detailed QAOA Circuit Description
Figure 15 gives an explicit description of the QAOA circuit formulation that is used in this study to sample this specific class of current IBM-processor defined two-qubit gate interaction graph. This QAOA circuit description can be extended to arbitrary $p$ by repeating this same $p=1$ structure. All QAOA simulations used the standard X mixer. All qubits are measured at the end of the simulation in the computational Z-basis.
@@
Appendix B LOWESA Simulations
@@
Appendix C PEPS Tensor Network Contraction Methods
@@
In Figure 17 we show comparisons using these different approximate tensor network contraction methods. Here we show examples of PEPS expectation values on some subset of the QAOA angle simulations on the $156$-qubit hardware-compatible Ising model instance, some of which show absolute agreement among the three tensor network contraction methods (in particular at $p=1$) and others which show small disagreement between expectation values. In general we see high agreement between boundary MPS, CTM, and BP methods, and therefore we use BP for all PEPS QAOA simulations reported in this study (with the exception of a handful of examples, such as in Figure 17, where the agreement is not absolutely 0), because it is more computationally efficient than the other three methods.
Figure 18: PEPS convergence plots for $156$-qubit QAOA circuits, using the other 88 trained angle sets from $16$-qubit instances, continuing from Figure 12.
Appendix D Remaining PEPS Convergence Plots
Figure 18 plots a complete set of PEPS convergence results for the remaining $8$ distinct sets of “QAOA schedules” trained on $16$-qubit problem instances. Notably, the convergence properties of each set of QAOA angles, shown in the form of different source training problem instances in each sub-plot, vary considerably. Ranging from a lack of strong convergence at high p, such as on the 16Q-8 instance, to very strong convergence in the case of the all-negative coefficient Ising model case (suggesting a lack of high entanglement produced by that set of QAOA circuits).
All of the convergence plots in Figure 18, as well as Figure 12, are normalized to use the same log-scale y-axis in the interval from 99 to $10^{-6}$.

### v2

@@ III.4 Validating Large Scale QAOA Parameter Transfer with PEPS Simulations
Fig. 8 plots the Hamiltonian expectation value as a function of $p$, for the complete set of $12$ trained sets of QAOA angles, using PEPS with a bond dimension of $D=64$, for the $127$, $133$, and $156$ qubit Ising model instances. The goal here is to validate whether the parameter transfer continues to work at larger problem sizes. Indeed, we see that for some of the QAOA angles, parameter transfer works remarkably well – allowing convergence to nearly the true ground-state. However, similarly to the smaller scale statevector simulations of Figs. 4, 14 and 15, we see that for some sets of parameters, there exist non-monotonic fluctuations, dips, and in one case there is a single set of QAOA angles which results in a significantly worse performance (note that this angle is shown explicitly as the highly fluctuating schedule shown in the 5th row from the top of Fig. 1). This result shows a single case where non-smooth QAOA angles can perform well on the small test case problem instance, but then fail to perform well when those parameter are transferred up to larger problem sizes – we hypothesize that this is likely a rare artifact associated with having converged to close the ground-state energy, although additional investigation is required on this point.
In summary, Fig. 8 shows that QAOA parameter transfer can work well across instances which have a parameter-transfer Ising model size difference (between the trained instance and the tested instance) of up to $140$ discrete variables, and in this case allows us to get near to the true ground-state, assuming a noise-less quantum computation (approximated by PEPS with a bond dimension of $D=64$).
Figure 9: Validating the QAOA angle quality, trained on the set of $12$ distinct $27$-qubit Ising model instances, applied to the $156$-qubit (left), $133$-qubit (middle), and $127$-qubit (right) instances, using $D=64$ PEPS simulations with belief propagation for approximate tensor network contraction. Ground state energy not marked because these expectation values $\leq 7$ do not reach near the ground state, as opposed to Fig. 8 where the mean expectation value does approach the true ground state energy.
@@
Figure 11: Ground-state energy QAOA sampling rate (probability of sampling the ground state) from $D=64$ and $D=32$ PEPS simulations, across the ensemble of $16$-qubit trained angles, on the $156$-qubit Ising model (left), $133$-qubit Ising model (middle) and the $127$-qubit Ising model (right). Interestingly, the $133$-qubit instance ground-state is rarely sampled, in comparison to the other two models. The ground-state sampling proportion is estimated using $1000$ samples, per parameter, (10,00010{,}000 in the case of the $133$-qubit case, to reduce the effect of shot noise due to the low ground-state energy sampling rate) and each sub-plot y-axis is different for each Ising model.
Fig. 9 shows similar trends of monotonic expectation value improvement as a function of $p$ from the $27$-qubit instance trained angles, as we would want from good QAOA evolution – but with one exception of the entirely positive coefficient Ising model having a non-monotonic dip in performance. Comparing Fig. 8 and Fig. 9, there is little difference between the angles that were trained either the slightly smaller ($16$ qubit) or larger ($27$ qubit) instances.
@@
Interestingly, one of the primary conclusions from all of these QAOA parameter transfer calculations is that overall, coefficient choice does not significantly impact the transfer of parameters (the only clear exception to this is that parameters from the positive coefficient case do not monotonically perform well on the other instances). Even very extreme coefficient choices, and random coefficient choices, at sufficiently large $p$ tend to work well. This suggests that these parameters could work well because the Hamiltonian connectivity structure was the same for all instances, however we do not evaluate parameter transfer across different connectivity graphs.
@@
Fig. 10 plots the relative expectation value differences from PEPS simulations at different bond dimensions run on the largest Ising model we consider in this study ($156$ spins). This shows that the tensor network simulations clearly converge at small $p$ (typically $p\leq 7$) and typically fail to converge as strongly at larger $p$, with slower convergence occurring in the intermediate values of $p$ typically $p=10$ to =20=20. Moreover, the trend of PEPS convergence is quite distinct for each set of fixed QAOA angles. This empirical convergence evidence reflects what we would expect from highly optimized QAOA evolution [99], similarly to quantum annealing, where in the middle of the evolution the system has high entanglement entropy, thus making it harder for the PEPS simulation to converge.
@@
Fig. 18 in Appendix E shows the convergence properties of the remaining set of trained QAOA angles, run on the single $156$-qubit instance. Interestingly, this lack of full convergence at intermediate p values (typically this is approximately between $p=10$ and $p=30$) suggests that those specific QAOA angles result in circuits which are harder to approximate classically using tensor network methods. This convergence trend manifests across most of the angles, with the one exception of the all negative-coefficient angle set shown in Fig. 18. However, the larger $p$ simulations, which use slightly different QAOA angles, are then closer to convergence. This suggests that there do exist particular QAOA parameter combinations for these circuits which are hard to classically simulate, however, unexpectedly the optimized angles, which typically approach the true ground-state energy, are easier to approximate classically. This could be related to the properties of the ground-states themselves which are approached in the high-$p$ regime, which can be expressed as a linear combination of many product states (i.e., due to high degeneracy of the Ising models).
@@
Convergence of the PEPS simulations would resemble a clear progression downward on the y-axis of logarithm expectation value difference in Fig. 10. Eventually at very large bond dimension, convergence results in the difference between bond dimensions (which are increasing as powers of 2) which have zero difference between each other. This effect is really only seen very strongly at low-p depth in Fig. 10, and in some cases we see a strong lack of convergence at sufficiently high p, indicating those expectation values have higher uncertainty, even at $D=128$.
@@
Lastly, a natural question in regards to quantum computation is the effect of sampling – for example Fig. 8 showed only the mean expectation value from the PEPS simulation, namely because all quantum computation requires measuring the state of the qubits at the end of the computation. Fig. 11 shows sampling from the PEPS simulation at $D=64$ using $1000$ samples (for each QAOA angle, at each $p$), where we quantify the sampled ground-state energy proportion. Interestingly, this shows that some of the parameter transfer angles work very well for ground-state sampling, typically well exceeding a proportion of $0.1$, while others do not perform well. This suggests in general for parameter transfer, using an ensemble of distinct QAOA angles is a good technique. However, the $133$-qubit instance ground-state energy is very rarely sampled. The gap between the best-performing angle and the ground-state energy of the $133$-qubit in Fig. 8 is larger than the gap for the other two instances, and therefore the low ground-state sampling rate in Fig. 11 is consistent with the noiseless expectation values. This suggests that the low ground state (GS) sampling rate is an instance-dependent result, and the performance of these fixed angles likely has high variability across different random-coefficient problem instances. Fig. 11 reports only the QAOA angles originating from the training on $16$-qubit instances since the $27$-qubit instances were only trained up to $p=7$, and in that QAOA depth regime we verified with PEPS simulations that GS energy sampling rate (using $1000$ samples per $p$) is zero.
@@ III.5 Validating QAOA PEPS Simulations with MPS Simulations
The MPS simulations are performed with $\chi=2^{k}$, for $k=7,\ldots,11$ and shown in Fig. 12 for the $156$ qubit QAOA circuits. Panels (a) and (b) display the error in the energy $\Delta E$ as a function of $p$ for several values of MPS bond dimension $\chi$. Here, the error $\Delta E$ quantifies the difference between our most accurate PEPS energy approximation and the MPS approximation computed with a given $\chi$. It is defined as:
@@ IV Discussion and Conclusion
However, parameter transfer does have limitations. Concretely, the following limitations exist for QAOA parameter transfer, primarily for the task of training on smaller problem instances and applying those parameters to larger problems. The most direct limitation is that once the QAOA training saturates the maximum approximation ratio of $1$, meaning no further improvement is possible, then the training can not supply good angles past that saturation point (for larger $p$) to use on any larger problem instances. This saturation is an issue specifically when we generally expect higher QAOA depth simulations to be required, to reach a certain threshold of accuracy (such as a ground-state sampling rate), as problem size increases. This limit necessitates the development of other techniques to be able to efficiently compute good QAOA angles, such as extrapolating to higher QAOA depths based on the already learned angles.
@@ Appendix A Detailed QAOA Circuit Description
Fig. 13 gives an explicit description of the QAOA circuit formulation that is used in this study to sample this specific class of current IBM-processor defined two-qubit gate interaction graph. This QAOA circuit description can be extended to arbitrary $p$ by repeating this same $p=1$ structure. All QAOA simulations used the standard X mixer. All qubits are measured at the end of the simulation in the computational Z-basis.
@@
Appendix B All Positive and Negative Donor QAOA Angles Parameter Transfer on $27$ and $16$ Qubit Instances
Fig. 14 shows parameter transfer performance for another set of fixed trained QAOA angles, specifically trained on the $16$-qubit Ising model with only negative coefficients. Fig. 15 shows the same thing but for the $16$-qubit Ising model with all positive coefficients.
Figs. 4 and 14 and 15 all show that the different trained QAOA angles do transfer onto other problem instances. In particular at sufficiently large $p$, these parameters do eventually show continuing improvement in energy as a function of $p$. The general finding is that despite very different random coefficient $\{+1,-1\}$ distributions, the QAOA angles still can be effectively transferred.
Appendix C LOWESA Numerical Simulations
@@
Appendix D PEPS Tensor Network Contraction Methods
@@
In Fig. 17 we show comparisons using these different approximate tensor network contraction methods. Here we show examples of PEPS expectation values on some subset of the QAOA angle simulations on the $156$-qubit hardware-compatible Ising model instance, some of which show absolute agreement among the three tensor network contraction methods (in particular at $p=1$) and others which show small disagreement between expectation values. In general we see high agreement between boundary MPS, CTM, and BP methods, and therefore we use BP for all PEPS QAOA simulations reported in this study (with the exception of a handful of examples, such as in Fig. 17, where the agreement is not absolutely 0), because it is more computationally efficient than the other three methods.
Figure 18: PEPS convergence plots for $156$-qubit QAOA circuits, using the other 88 trained angle sets from $16$-qubit instances, continuing from Fig. 10.
Appendix E Remaining PEPS Convergence Plots
Fig. 18 plots a complete set of PEPS convergence results for the remaining $8$ distinct sets of “QAOA schedules” trained on $16$-qubit problem instances. Notably, the convergence properties of each set of QAOA angles, shown in the form of different source training problem instances in each sub-plot, vary considerably. Ranging from a lack of strong convergence at high p, such as on the 16Q-8 instance, to very strong convergence in the case of the all-negative coefficient Ising model case (suggesting a lack of high entanglement produced by that set of QAOA circuits).
All of the convergence plots in Fig. 18, as well as Fig. 10, are normalized to use the same log-scale y-axis in the interval from 99 to $10^{-6}$.